# A Metamorphic Testing Perspective on Knowledge Distillation for Language Models of Code: Does the Student Deeply Mimic the Teacher?

### v1

@@ 1 Introduction
Transformer-based vaswani2017attention language models of code have achieved state-of-the-art performance across a wide range of software analytics tasks, including clone detection arshad2022codebert, code summarization ahmed2024automatic, vulnerability prediction ding2024vulnerability, and code search chen2024code. However, despite their impressive capabilities, the adoption of these models remains limited, particularly in consumer-grade devices (e.g., laptops) and latency-sensitive applications, due to their substantial computational demands, slow inference speed, and considerable energy consumption and carbon footprint shi2022compressing, schwartz2020green. For instance, CodeBERT, a widely used code language model with 125 million parameters and a size of 476 MB, can incur up to 1.5 seconds of latency, consume 0.32 kWh of energy, and emit 0.14 kg of CO2 on consumer-grade laptops. [^1] When integrated into IDEs, CodeBERT may be invoked thousands of times daily by a developer, reflecting typical usage patterns hellendoorn2019code. To address these practical challenges, the software engineering research community has recently begun exploring knowledge distillation methods hinton2015distilling to develop extremely compact models that reduce model size, lower inference latency, and mitigate environmental impact while maintaining task performance shi2022compressing, shi2024greening, shi2024efficient, wei2023towards, saad2024alpine, d2024compression, chen2025smaller.
@@
Knowledge distillation (KD) transfers knowledge from a large teacher model to a smaller student [^2] In this study, the term student model refers to the compressed model obtained through knowledge distillation, while the term teacher model denotes the original uncompressed model. model, maintaining comparable accuracy while requiring significantly fewer computational resources. In this context, Shi et al. shi2024greening introduced Compressor, a genetic algorithm-based approach that compresses code-specialized language models (e.g., CodeBERT feng2020codebert) from 476 MB and 125 million parameters to only 3 MB with minimal loss in accuracy. Although KD effectively reduces computational costs and environmental impact, student model performance is typically evaluated using traditional accuracy-based metrics that assess whether the student model’s task performance on ground-truth labels remains comparable to that of the teacher model. However, this form of evaluation offers only a surface-level view and may overlook critical discrepancies in behavioral fidelitythe extent to which a student model deeply mimics the predictive behavior and internal representations of its teacher counterpart.
@@
We investigate the student model’s ability to mimic the teacher by performing simple yet effective adversarial attacks yang2022natural, zhang2020generating, zeng2022extensive that rename identifiers based on the contextual semantics of each code snippet. The empirical results reveal that although the student models achieve performance comparable to the teacher before adversarial perturbations, their performance degrades substantially under adversarial attacks, with up to 285% greater performance loss than the teacher model. This finding indicates that the student model does not deeply mimic the teacher’s predictive behavior and internal representations, highlighting the limitations of the traditional accuracy-based evaluation, which often overlooks more profound behavioral discrepancies between teacher and student models. It also underscores the need for a nuanced evaluation framework that goes beyond accuracy to assess how deeply a student model mimics the predictive behavior and internal representations of its teacher, providing a more comprehensive understanding of model alignment under knowledge distillation.
@@
To fill this gap, we introduce MetaCompress, a testing framework inspired by metamorphic testing principles in software engineering chen2018metamorphic, to evaluate the behavioral fidelity of the student model in knowledge distillation. In contrast to traditional accuracy-based evaluation, which compares model performance against ground-truth labels, MetaCompress performs a head-to-head comparison of student and teacher model outputs across a suite of behavior-preserving metamorphic relations, enabling a more nuanced assessment of behavioral fidelity. We design four metamorphic relations, grounded in strong empirical evidence from manual investigations, that capture discrepancies in the behavioral fidelity of student models, ranging from prediction-level disagreement to divergences in the probability distribution.
@@
We empirically evaluate MetaCompress on two widely studied software engineering tasks, clone detection and vulnerability prediction, using two popular language models of code, CodeBERT feng2020codebert and GraphCodeBERT guo2020graphcodebert. To distill their smaller student models, we apply three state-of-the-art knowledge distillation techniques designed for language models of code: Compressor shi2022compressing, AVATAR shi2024greening, and MORPH panichella2025metamorphic. Our extensive experiments show that MetaCompress can reveal behavioral fidelity discrepancies of up to 62% across four metamorphic relations that traditional accuracy-based metrics fail to detect.
@@ 2.1 Language Models of Code
Transformer-based vaswani2017attention language models of code have demonstrated remarkable performance across a wide range of software analytics tasks, such as vulnerability prediction, clone detection, code summarization, code completion, code comment generation, and code search feng2020codebert, guo2020graphcodebert, lu2021codexglue, ahmad2021unified, wang2021codet5, devlin2018bert, liu2019roberta, sanh2019distilbert, husain2019codesearchnet, hou2024large. These models learn general language representations from large amounts of unlabeled data through a two-step process: (i) pre-training and (ii) fine-tuning. During pre-training, the model employs a self-supervised objective to acquire a general understanding of source code structures and patterns. Once the training is complete, the model can be fine-tuned by updating a few parameters for specific tasks with smaller datasets, making it more efficient in terms of time and computational resources compared to training a model from scratch devlin2018bert.
Language models for code can be broadly classified into three categories: (1) encoder-only, (2) decoder-only, and (3) encoder-decoder models zeng2022extensive. Encoder-only models (e.g., CodeBERT feng2020codebert and GraphCodeBERT guo2020graphcodebert) process the entire input simultaneously and are primarily used for discriminative tasks, such as classification (e.g., vulnerability prediction and clone detection). Decoder-only models (e.g., CodeGPT lu2021codexglue, StarCoder lozhkov2024starcoder, and Codex chen2021evaluating) generate output token by token and are typically employed in generative tasks, such as code generation, code completion, and summarization. Encoder-decoder models (e.g., PLBART ahmad2021unified and CodeT5 wang2021codet5) encode the input into a latent representation and subsequently decode it into an output sequence, making them suitable for translation-style tasks such as code summarization and programming language translation. Among these categories, encoder only models, particularly CodeBERT and GraphCodeBERT, have been widely and effectively applied to various downstream software analytics tasks and model compression studies shi2022compressing, shi2024greening, shi2024efficient, panichella2025metamorphic, liu2026pioneer, as evidenced by prior empirical evaluations lu2021codexglue. Consequently, these models serve as the primary focus of our study.
@@ 2.2 Model Compression
A significant barrier to the real-world deployment of language models of code is their substantial computational overhead, high memory demands, and environmental impact (e.g., carbon emissions) schwartz2020green, xu2023survey. To mitigate these challenges and advance sustainability, a widely adopted strategy known as model compression offers a solution that reduces the resource consumption of such models without significantly compromising performance. Model compression refers to the process of reducing a large, computationally intensive model into a smaller, more efficient version suitable for deployment on resource-constrained devices (e.g., consumer-grade laptops). In addition to deployment benefits, model compression improves inference speed and enhances overall resource efficiency. Algorithmically, the fundamental model compression techniques include: pruning sanh2020movement, quantization zafrir2019q8bert, knowledge distillation hinton2015distilling, neural architecture search elsken2019neural, weight sharing & hashing chen2015compressing, and low-rank factorization sainath2013low-each method uniquely contributes to model optimization. Among these, pruning, quantization, and knowledge distillation (KD) have emerged as the most prominent and widely adopted approaches for compressing language models of code in recent software engineering research. shi2022compressing, shi2024greening, wei2023towards, d2024compression, chen2025smaller, liu2026pioneer, afrin2025quantization.
@@ 2.3 Metamorphic Testing
Metamorphic testing (MT) is a property-based testing paradigm widely adopted for validating software systems in the absence of a test oracle or when one is difficult to define, a challenge commonly referred to as the oracle problem chen2018metamorphic, zhang2014search. Rather than verifying the correctness of individual outputs in isolation, MT addresses this issue by evaluating whether the program under test exhibits expected relationships between outputs across a set of well-defined metamorphic relations (MRs). An MR specifies a necessary property that captures the expected relationship between multiple inputs and their corresponding outputs. These relations are derived from the problem domain’s inherent semantics and define how the output should change (or remain consistent) when the input is systematically modified. This paradigm is beneficial for testing complex systems such as machine learning models, scientific computing software, or simulation tools, where correctness is difficult to assert for individual outputs but consistency across related cases can be meaningfully assessed yang2025hallucination, xie2011testing, chen2020metamorphic, xiao2022metamorphic, segura2018metamorphic.
@@ 2.4 Adversarial Attack
An adversarial attack refers to the deliberate process of crafting adversarial examples by introducing subtle perturbations to the original inputs. Although these perturbed instances appear almost identical to their original counterparts from a human perspective, they can cause the model to yield erroneous or inconsistent outputs. The concept of adversarial attack was first introduced by Zegedy et al. szegedy2013intriguing, who demonstrated that state-of-the-art image classifiers can be misled by carefully designed pixel-level modifications that remain imperceptible to human observers. Formally, given a classifier $f:X\rightarrow Y$ that correctly predicts the label $y_{truth}\in Y$ for an input $x\in X$, an adversarial attack generates a perturbed version $x^{\prime}$ such that $f(x^{\prime})\neq f(x)=y_{truth}$.
In continuous domains such as images, producing visually indistinguishable adversarial examples is relatively straightforward. However, generating adversarial examples for discrete domains like source code is considerably more complex due to the rigid syntactic and grammatical structure of programming languages zhang2022towards, zhang2020generating. As a result, adversarial examples designed to attack language models of code must meet three fundamental requirements. First, the perturbation must effectively mislead the model and reduce its predictive performance. Second, the perturbed code must remain syntactically correct, adhering to the formal rules of the programming language; for example, in Python, identifiers cannot begin with digits and may only contain letters, digits, or underscores. Third, the perturbed code $x^{\prime}$ must preserve the semantics of the original code $x$, maintaining identical functionality and behavior when executed under the same conditions. When an input instance is modified in accordance with these constraints to generate a perturbed version, the resulting instance constitutes an adversarial example; if it causes the model to produce an incorrect or inconsistent output, it is regarded as an adversarial attack.
@@ 2.5 Adversarial Robustness
Adversarial robustness describes a model’s capacity to maintain its predictive performance and stability when exposed to adversarial perturbations. A model is considered adversarially robust if it continues to generate consistent, accurate outputs despite deliberate input modifications intended to mislead it. In the context of machine learning, this concept reflects a model’s resilience to adversarial attacks that exploit its sensitivity to small perturbations in the input space szegedy2013intriguing, zhang2022towards.
@@ 2.6 Kullback–Leibler Divergence
The Kullback-Leibler (KL) divergence kullback1951information is a measure of how one probability distribution $Q$ diverges from a reference probability distribution $P$. For discrete distributions defined over the same support, it is mathematically expressed as:
@@ 3.1.1 Knowledge Distillation Approaches
In recent years, several knowledge distillation (KD) approaches have been proposed for language models of code, including Compressor shi2022compressing, AVATAR shi2024greening, SODA chen2025smaller, MORPH panichella2025metamorphic, and PIONEER liu2026pioneer. In this study, we select Compressor, AVATAR, and MORPH as our KD approaches to generate student models, as the replication package for SODA is not fully available [^4] https://shorturl.at/Se0f7. Furthermore, both MORPH and PIONEER incorporate robustness as an additional objective of knowledge distillation; however, because MORPH optimizes a broader set of objectives than PIONEER, we exclude PIONEER from our empirical investigation.
@@
In “(2)”, $p_{i}$ and $q_{i}$ represent the teacher and student model logits, respectively, with $T$ denoting the temperature parameter controlling the softmax distribution hinton2015distilling. While the teacher logits remain fixed, the student logits are updated throughout the training process. It is essential to note that we utilize off-the-shelf compressed versions of CodeBERT and GraphCodeBERT, produced by MORPH and AVATAR, for clone detection and vulnerability prediction tasks in this empirical evaluation.
@@ 3.1.2 Datasets and Subjects
To minimize experimental bias in evaluating whether the student model deeply mimics the teacher’s predictive behavior and internal representations during knowledge distillation, we adopt the same models and datasets used in the knowledge distillation–based approaches described in Section 3.1.1. Specifically, we select CodeBERT feng2020codebert and GraphCodeBERT guo2020graphcodebert models and two well-studied downstream tasks: clone detection and vulnerability prediction, which serve as standard benchmarks in knowledge distillation research for software engineering shi2022compressing, shi2024greening, shi2024efficient, panichella2025metamorphic, liu2026pioneer.
Clone detection: (CD) This task aims to determine whether two given code snippets are syntactically or semantically equivalent. To streamline the fine-tuning process, we adopt the filtered BigCloneBench dataset svajlenko2014towards provided by Wang et al. wang2020detecting. The final dataset includes 90,102 training examples, along with 4,000 examples each for validation and testing. Vulnerability prediction: (VP) This task aims to determine whether a given code snippet contains a vulnerability. We conduct our experiments using the Devign [^5] https://sites.google.com/view/devign dataset, introduced by Zhou et al. zhou2019devign. The dataset comprises 27,318 functions sourced from two widely used open-source C libraries, FFmpeg and Qemu. Each function is manually annotated to indicate whether it contains vulnerabilities. We use the CodeXGLUE benchmark splits lu2021codexglue, including training, development, and test sets with 27,318, 2,732, and 2,732 examples, respectively.
CodeBERT is a pre-trained language model that shares the same architecture and a similar pre-training strategy as RoBERTa liu2019roberta. It consists of twelve Transformer layers, each with 12-headed multi-head attention mechanisms and a hidden size of 768. CodeBERT is pre-trained using two objectives: masked language modeling, which predicts masked tokens, and replaced token detection, which identifies substituted tokens in the input. The model is trained on bimodal data comprising natural language–code pairs across six programming languages from the CodeSearchNet dataset husain2019codesearchnet. GraphCodeBERT shares CodeBERT’s architecture but incorporates data-flow information from source code during pre-training. Additionally, it introduces a pre-training task that predicts whether a data flow edge exists between two nodes.

### v2

@@ 1 Introduction
Transformer-based Vaswani et al. [2017] language models of code have achieved state-of-the-art performance across a wide range of software analytics tasks, including clone detection Arshad et al. [2022], code summarization Ahmed et al. [2024], vulnerability prediction Ding et al. [2024], and code search Chen et al. [2024]. However, despite their impressive capabilities, the adoption of these models remains limited, particularly in consumer-grade devices (e.g., laptops) and latency-sensitive applications, due to their substantial computational demands, slow inference speed, and considerable energy consumption and carbon footprint Shi et al. [2023], Schwartz et al. [2020]. For instance, CodeBERT, a widely used code language model with 125 million parameters and a size of 476 MB, can incur up to 1.5 seconds of latency, consume 0.32 kWh of energy, and emit 0.14 kg of CO2 on consumer-grade laptops. [^1] When integrated into IDEs, CodeBERT may be invoked thousands of times daily by a developer, reflecting typical usage patterns Hellendoorn et al. [2019]. To address these practical challenges, the software engineering research community has recently begun exploring knowledge distillation methods Hinton et al. [2015] to develop extremely compact models that reduce model size, lower inference latency, and mitigate environmental impact while maintaining task performance Shi et al. [2023, 2024a, 2024b], Wei et al. [2023], Saad et al. [2024], d’Aloisio et al. [2024], Chen et al. [2025].
@@
Knowledge distillation (KD) transfers knowledge from a large teacher model to a smaller student [^2] In this study, the term student model refers to the compressed model obtained through knowledge distillation, while the term teacher model denotes the original uncompressed model. model, maintaining comparable accuracy while requiring significantly fewer computational resources. In this context, Shi et al. Shi et al. [2024a] introduced Compressor, a genetic algorithm-based approach that compresses code-specialized language models (e.g., CodeBERT Feng et al. [2020]) from 476 MB and 125 million parameters to only 3 MB with minimal loss in accuracy. Although KD effectively reduces computational costs and environmental impact, student model performance is typically evaluated using traditional accuracy-based metrics that assess whether the student model’s task performance on ground-truth labels remains comparable to that of the teacher model. However, this form of evaluation offers only a surface-level view and may overlook critical discrepancies in behavioral fidelitythe extent to which a student model deeply mimics the predictive behavior and internal representations of its teacher counterpart.
@@
We investigate the student model’s ability to mimic the teacher by performing simple yet effective adversarial attacks Yang et al. [2022], Zhang et al. [2020a], Zeng et al. [2022] that rename identifiers based on the contextual semantics of each code snippet. The empirical results reveal that although the student models achieve performance comparable to the teacher before adversarial perturbations, their performance degrades substantially under adversarial attacks, with up to 285% greater performance loss than the teacher model. This finding indicates that the student model does not deeply mimic the teacher’s predictive behavior and internal representations, highlighting the limitations of the traditional accuracy-based evaluation, which often overlooks more profound behavioral discrepancies between teacher and student models. It also underscores the need for a nuanced evaluation framework that goes beyond accuracy to assess how deeply a student model mimics the predictive behavior and internal representations of its teacher, providing a more comprehensive understanding of model alignment under knowledge distillation.
@@
To fill this gap, we introduce MetaCompress, a testing framework inspired by metamorphic testing principles in software engineering Chen et al. [2018], to evaluate the behavioral fidelity of the student model in knowledge distillation. In contrast to traditional accuracy-based evaluation, which compares model performance against ground-truth labels, MetaCompress performs a head-to-head comparison of student and teacher model outputs across a set of behavior-preserving metamorphic relations, enabling a more nuanced assessment of behavioral fidelity. We design four metamorphic relations, grounded in strong empirical evidence from manual investigations, that capture discrepancies in the behavioral fidelity of student models, ranging from prediction-level disagreement to divergences in the probability distribution.
@@
We empirically evaluate MetaCompress on two widely studied software engineering tasks, clone detection and vulnerability prediction, using two popular language models of code, CodeBERT Feng et al. [2020] and GraphCodeBERT Guo et al. [2020]. To distill their smaller student models, we apply three state-of-the-art knowledge distillation techniques designed for language models of code: Compressor Shi et al. [2023], AVATAR Shi et al. [2024a], and MORPH Panichella [2025]. Our extensive experiments show that MetaCompress can reveal behavioral fidelity discrepancies of up to 62% across four metamorphic relations that traditional accuracy-based metrics fail to detect.
@@ 2.1 Language Models of Code
Transformer-based Vaswani et al. [2017] language models of code have demonstrated remarkable performance across a wide range of software analytics tasks, such as vulnerability prediction, clone detection, code summarization, code completion, code comment generation, and code search Feng et al. [2020], Guo et al. [2020], Lu et al. [2021], Ahmad et al. [2021], Wang et al. [2021], Devlin [2018], Liu [2019], Sanh [2019], Husain et al. [2019], Hou et al. [2024]. These models learn general language representations from large amounts of unlabeled data through a two-step process: (i) pre-training and (ii) fine-tuning. During pre-training, the model employs a self-supervised objective to acquire a general understanding of source code structures and patterns. Once the training is complete, the model can be fine-tuned by updating a few parameters for specific tasks with smaller datasets, making it more efficient in terms of time and computational resources compared to training a model from scratch Devlin [2018].
Language models for code can be broadly classified into three categories: (1) encoder-only, (2) decoder-only, and (3) encoder-decoder models Zeng et al. [2022]. Encoder-only models (e.g., CodeBERT Feng et al. [2020] and GraphCodeBERT Guo et al. [2020]) process the entire input simultaneously and are primarily used for discriminative tasks, such as classification (e.g., vulnerability prediction and clone detection). Decoder-only models (e.g., CodeGPT Lu et al. [2021], StarCoder Lozhkov et al. [2024], and Codex Chen et al. [2021]) generate output token by token and are typically employed in generative tasks, such as code generation, code completion, and summarization. Encoder-decoder models (e.g., PLBART Ahmad et al. [2021] and CodeT5 Wang et al. [2021]) encode the input into a latent representation and subsequently decode it into an output sequence, making them suitable for translation-style tasks such as code summarization and programming language translation. Among these categories, encoder-only models, particularly CodeBERT and GraphCodeBERT, have been widely and effectively applied to various downstream software analytics tasks and model compression studies Shi et al. [2023, 2024a, 2024b], Panichella [2025], Liu et al. [2026], as evidenced by prior empirical evaluations Lu et al. [2021]. Consequently, these models serve as the primary focus of our study.
@@ 2.2 Model Compression
A significant barrier to the real-world deployment of language models of code is their substantial computational overhead, high memory demands, and environmental impact (e.g., carbon emissions) Schwartz et al. [2020], Xu and McAuley [2023]. To mitigate these challenges and advance sustainability, a widely adopted strategy known as model compression offers a solution that reduces the resource consumption of such models without significantly compromising performance. Model compression refers to the process of reducing a large, computationally intensive model into a smaller, more efficient version suitable for deployment on resource-constrained devices (e.g., consumer-grade laptops). In addition to deployment benefits, model compression improves inference speed and enhances overall resource efficiency. Algorithmically, the fundamental model compression techniques include: pruning Sanh et al. [2020], quantization Zafrir et al. [2019], knowledge distillation Hinton et al. [2015], neural architecture search Elsken et al. [2019], weight sharing & hashing Chen et al. [2015], and low-rank factorization Sainath et al. [2013]-each method uniquely contributes to model optimization. Among these, pruning, quantization, and knowledge distillation (KD) have emerged as the most prominent and widely adopted approaches for compressing language models of code in recent software engineering research. Shi et al. [2023, 2024a], Wei et al. [2023], d’Aloisio et al. [2024], Chen et al. [2025], Liu et al. [2026], Afrin et al. [2025].
@@ 2.3 Metamorphic Testing
Metamorphic testing (MT) is a property-based testing paradigm widely adopted for validating software systems in the absence of a test oracle or when one is difficult to define, a challenge commonly referred to as the oracle problem Chen et al. [2018], Zhang et al. [2014]. Rather than verifying the correctness of individual outputs in isolation, MT addresses this issue by evaluating whether the program under test exhibits expected relationships between outputs across a set of well-defined metamorphic relations (MRs). An MR specifies a necessary property that captures the expected relationship between multiple inputs and their corresponding outputs. These relations are derived from the problem domain’s inherent semantics and define how the output should change (or remain consistent) when the input is systematically modified. This paradigm is beneficial for testing complex systems such as machine learning models, scientific computing software, or simulation tools, where correctness is difficult to assert for individual outputs but consistency across related cases can be meaningfully assessed Yang et al. [2025], Xie et al. [2011], Chen et al. [2020], Xiao et al. [2022], Segura et al. [2018].
@@ 2.4 Adversarial Attack
An adversarial attack refers to the deliberate process of crafting adversarial examples by introducing subtle perturbations to the original inputs. Although these perturbed instances appear almost identical to their original counterparts from a human perspective, they can cause the model to yield erroneous or inconsistent outputs. The concept of adversarial attack was first introduced by Zegedy et al. Szegedy et al. [2013], who demonstrated that state-of-the-art image classifiers can be misled by carefully designed pixel-level modifications that remain imperceptible to human observers. Formally, given a classifier $f:X\rightarrow Y$ that correctly predicts the label $y_{truth}\in Y$ for an input $x\in X$, an adversarial attack generates a perturbed version $x^{\prime}$ such that $f(x^{\prime})\neq f(x)=y_{truth}$.
In continuous domains such as images, producing visually indistinguishable adversarial examples is relatively straightforward. However, generating adversarial examples for discrete domains like source code is considerably more complex due to the rigid syntactic and grammatical structure of programming languages Zhang et al. [2022, 2020a]. As a result, adversarial examples designed to attack language models of code must meet three fundamental requirements. First, the perturbation must effectively mislead the model and reduce its predictive performance. Second, the perturbed code must remain syntactically correct, adhering to the formal rules of the programming language; for example, in Python, identifiers cannot begin with digits and may only contain letters, digits, or underscores. Third, the perturbed code $x^{\prime}$ must preserve the semantics of the original code $x$, maintaining identical functionality and behavior when executed under the same conditions. When an input instance is modified in accordance with these constraints to generate a perturbed version, the resulting instance constitutes an adversarial example; if it causes the model to produce an incorrect or inconsistent output, it is regarded as an adversarial attack.
@@ 2.5 Adversarial Robustness
Adversarial robustness describes a model’s capacity to maintain its predictive performance and stability when exposed to adversarial perturbations. A model is considered adversarially robust if it continues to generate consistent, accurate outputs despite deliberate input modifications intended to mislead it. In the context of machine learning, this concept reflects a model’s resilience to adversarial attacks that exploit its sensitivity to small perturbations in the input space Szegedy et al. [2013], Zhang et al. [2022].
@@ 2.6 Kullback–Leibler Divergence
The Kullback-Leibler (KL) divergence Kullback and Leibler [1951] is a measure of how one probability distribution $Q$ diverges from a reference probability distribution $P$. For discrete distributions defined over the same support, it is mathematically expressed as:
@@ 3.1.1 Knowledge Distillation Approaches
In recent years, several knowledge distillation (KD) approaches have been proposed for language models of code, including Compressor Shi et al. [2023], AVATAR Shi et al. [2024a], SODA Chen et al. [2025], MORPH Panichella [2025], and PIONEER Liu et al. [2026]. In this study, we select Compressor, AVATAR, and MORPH as our KD approaches to generate student models, as the replication package for SODA is not fully available [^4] https://shorturl.at/Se0f7. Furthermore, both MORPH and PIONEER incorporate robustness as an additional objective of knowledge distillation; however, because MORPH optimizes a broader set of objectives than PIONEER, we exclude PIONEER from our empirical investigation.
@@
In “(2)”, $p_{i}$ and $q_{i}$ represent the teacher and student model logits, respectively, with $T$ denoting the temperature parameter controlling the softmax distribution Hinton et al. [2015]. While the teacher logits remain fixed, the student logits are updated throughout the training process. It is essential to note that we utilize off-the-shelf compressed versions of CodeBERT and GraphCodeBERT, produced by MORPH and AVATAR, for clone detection and vulnerability prediction tasks in this empirical evaluation.
@@ 3.1.2 Datasets and Subjects
To minimize experimental bias in evaluating whether the student model deeply mimics the teacher’s predictive behavior and internal representations during knowledge distillation, we adopt the same models and datasets used in the knowledge distillation–based approaches described in Section 3.1.1. Specifically, we select CodeBERT Feng et al. [2020] and GraphCodeBERT Guo et al. [2020] models and two well-studied downstream tasks: clone detection and vulnerability prediction, which serve as standard benchmarks in knowledge distillation research for software engineering Shi et al. [2023, 2024a, 2024b], Panichella [2025], Liu et al. [2026].
Clone detection: (CD) This task aims to determine whether two given code snippets are syntactically or semantically equivalent. To streamline the fine-tuning process, we adopt the filtered BigCloneBench dataset Svajlenko et al. [2014] provided by Wang et al. Wang et al. [2020]. The final dataset includes 90,102 training examples, along with 4,000 examples each for validation and testing. Vulnerability prediction: (VP) This task aims to determine whether a given code snippet contains a vulnerability. We conduct our experiments using the Devign [^5] https://sites.google.com/view/devign dataset, introduced by Zhou et al. Zhou et al. [2019]. The dataset comprises 27,318 functions sourced from two widely used open-source C libraries, FFmpeg and Qemu. Each function is manually annotated to indicate whether it contains vulnerabilities. We use the CodeXGLUE benchmark splits Lu et al. [2021], including training, development, and test sets with 27,318, 2,732, and 2,732 examples, respectively.
CodeBERT is a pre-trained language model that shares the same architecture and a similar pre-training strategy as RoBERTa Liu [2019]. It consists of twelve Transformer layers, each with 12-headed multi-head attention mechanisms and a hidden size of 768. CodeBERT is pre-trained using two objectives: masked language modeling, which predicts masked tokens, and replaced token detection, which identifies substituted tokens in the input. The model is trained on bimodal data comprising natural language–code pairs across six programming languages from the CodeSearchNet dataset Husain et al. [2019]. GraphCodeBERT shares CodeBERT’s architecture but incorporates data-flow information from source code during pre-training. Additionally, it introduces a pre-training task that predicts whether a data flow edge exists between two nodes.